OpenStudy (isabel☺):

What is the laplace transform of 9t^(-9) ?

OpenStudy (inkyvoyd):

were you provided a table?

OpenStudy (isabel☺):

ahm i lost it haha. but i think: is this one?

OpenStudy (isabel☺):

@inkyvoyd

OpenStudy (inkyvoyd):

that's for the positive exponents... I'm not sure you can even use the gamma function to extrapolate the negative factorial here

OpenStudy (inkyvoyd):

try evaluating the integral by definition.

OpenStudy (isabel☺):

whew i forgot calculus i'm sorry can you help me remember that? i'm so sorry

OpenStudy (isabel☺):

the answer in my notes is -9/s^(-8) . i don't know how that happened :(

OpenStudy (inkyvoyd):

if you're taking ODEs and you forgot calculus you're screwed anyways lol

OpenStudy (isabel☺):

you're arrogant.

OpenStudy (inkyvoyd):

I mean, if you're taking physics and you forget calculus you're screwed anyways lol... I might be arrogant, but unfortunately for you pointing that out doesn't change the fact that if you forgot calculus you're screwed lol. Just put down the formula for laplace, and we'll evaluate the integral and see where it came from lol.

OpenStudy (inkyvoyd):

\(\Huge \int_0^\infty f(t)e^{-st}dt\) plug in f(t) and let's go.

OpenStudy (isabel☺):

\[\int\limits_{0}^{\infty} 9t ^{-9} e ^{-st}dt\] ? ?

OpenStudy (inkyvoyd):

yeah

OpenStudy (inkyvoyd):

of course the hard part now is evaluating that sucker, but i'm sure we'll find a way

OpenStudy (isabel☺):

?

OpenStudy (unklerhaukus):

but p is negative

OpenStudy (inkyvoyd):

yeah we have a big problem here because the gamma function isn't defined for negative integers

OpenStudy (apple):

You can evaluate that integral using integration by parts

OpenStudy (inkyvoyd):

You can't evaluate the integral that we posted lol. I checked already; it's a dead end.

OpenStudy (apple):

Yeah lol i tried it didn't work

OpenStudy (inkyvoyd):

It's because the integral doesn't actually exist lol http://math.stackexchange.com/questions/147126/laplace-transform-of-1-t

OpenStudy (inkyvoyd):

I don't know how your notebook has -9/s^(-8) becauese I'm pretty sure it's just wrong

OpenStudy (isabel☺):

oh, thanks . ya my prof is not really that good in math that's what he said. thanks for your time

OpenStudy (isabel☺):

wait what? inverse laplace? haha ok i get it now